OpenStudy (anonymous):

subtract 3 1/8 from 5

OpenStudy (anonymous):

have to show work

OpenStudy (anonymous):

I would change the 3 1/8 into an improper fraction. To do this, multiply 8•3 and add 1 and put that number over the denominator. So, 25/8. Next, change 5 into a fraction with 8 in the denominator. 40/8 works. 40/8 - 25/8 = 15/8

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

I don't know if you have to change the fraction into a mixed number or decimal...do your teachers have a preference?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well, mixed number = 1 7/8 and decimal = 1.875

OpenStudy (anonymous):

got it